OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos2x + 2 cos x + 1 = 0

OpenStudy (p0sitr0n):

is it cos(2x) or cos^2 x ?

OpenStudy (anonymous):

Second choice lol

OpenStudy (p0sitr0n):

say cosx = u u^2 + 2u +1 =0 (u+1)(u+1)=0 root : u = -1 cosx = -1 when x = pi root = pi

OpenStudy (anonymous):

\[x = \pi \]

OpenStudy (anonymous):

x = 2π x = π x = pi divided by four., seven pi divided by four. x = pi divided by two., three pi divided by two.

OpenStudy (p0sitr0n):

x = pi